OpenStudy (anonymous):

Help :) The fraction 1-1/a divided by 1-1/a^2 is equiv. to... Solve for x: 3/x-2+8/x^2-4=1

OpenStudy (anonymous):

For the first question try setting it up this way: (1 - 1/a) / (1 - 1/a^2) = (a - 1)/a / (a^2 - 1) / a^2 Instead of dividing, multiply by the reciprocal, then rewrite a^2 - 1 as (a + 1)(a - 1). You will find that most everything cancels. As for the second question, try multiplying everything by x^2. That way you will get 3x - 2x^2 + 8 - 4x^2 = x^2

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

:)